A boy determines the greatest strength of his hands by means of a forcématėsi. Find the work done by him, if the force shown by the forcemeter is 600 N. The coefficient of elasticity of the forcemeter system is k = 10 000 N / m.


1
Expert's answer
2021-04-13T06:38:26-0400

"W=\frac{kx^2}{2}"


"F=kx\to x=F/k"


So, we have


"W=\frac{k\cdot F^2}{2\cdot k^2}=\frac{F^2}{2k}=\frac{600^2}{2\cdot10000}=18\ (J)" . Answer